# kendallknight: An R Package for Efficient Implementation of Kendall’s Correlation Coefficient Computation

### v3

@@
m.sepulveda@mail.utoronto.ca
ORCID $0$-0003-1017-7574
Department of Political Science and Munk School of Global Affairs and Public Policy
University of Toronto
@@ 3 Definitions
An array that represents a binary tree has a search operation with a computational complexity of $O(\log(n))$ and an insertion operation with a computational complexity of $O(n)$ (Abrevaya 1999; Christensen 2005). The resulting computational complexity of the search and insert operations in an array is $O(n)$ (Emara 2024). Repeating thhe search and insert operation for each element in the array results in a computational complexity of $O(n^{2})$, resulting in the same computational complexity as the naive implementation.
@@ 5 Benchmarks
In order to avoid distorted results, we used the bench package to run the benchmarking tests in a clean R session without other processes running in the background (e.g., such as automatic updates) on a ThinkPad X1 Carbon with an Intel Core i7-1185G7 with eight cores running Pop!_OS 22.04.
@@ 7 Installation and Usage
It can also be installed from GitHub using the following command:
@@
## \$statistic
## [1] 0.1284848
## \$p_value
## [1] 0.970893
## \$alternative
## [1] "alternative hypothesis: true tau is less than 0"

### v4

@@
Department of Political Science, University of Toronto
Munk School of Global Affairs and Public Policy, University of Toronto
Corresponding author: m.sepulveda@mail.utoronto.ca
@@ 3 Definitions
An array that represents a binary tree has a search operation with a computational complexity of $O(\log(n))$ and an insertion operation with a computational complexity of $O(n)$ (Abrevaya 1999; Christensen 2005). The resulting computational complexity of the search and insert operations in an array is $O(n)$ (Emara 2024). Repeating the search and insert operation for each element in the array results in a computational complexity of $O(n^{2})$, resulting in the same computational complexity as the naive implementation.
@@ 5 Benchmarks
The benchmarks were conducted on a ThinkPad X1 Carbon Gen 9 with the following specifications:
  • Processor: Intel Core i7-1185G7 with eight cores
  • Memory: 16 GB LPDDR4Xx-4266
  • Operating System: Pop!_OS 22.04 based on Ubuntu 22.04
  • R Version: 4.4.1
  • BLAS Library: OpenBLAS 0.3.20
@@ 7 Installation and Usage
# GitHub
@@
\$statistic
[1] 0.1288889
\$p_value
[1] 0.971286
\$alternative
[1] "alternative hypothesis: true tau is less than 0"